Galactose and glucose compose lactose, which is referred to as milk sugar. What is the molecular formula of galactose if the empirical formula is CH2O, and the approximate molar mass is 180 g/mole
1
Expert's answer
2019-04-17T04:17:12-0400

Mmol = 180 g/mol.

Memp(CH2O) = 12 g/mol + 2 g/mol + 16 g/mol = 30 g/mol.

Mmol = Memp × n = 30 g/mol × n = 180 g/mol.

n = 180/30 = 6.

Therefore, the molecular formula of galactose: C6H12O6.